# UniMODE: Unified Monocular 3D Object Detection

### v4

@@ 3.3 Uneven BEV Grid
Therefore, the BEV grid sizes of current outdoor detectors are typically large to accommodate the vast perception range, while those of indoor detectors are small because of the intricate indoor scenes. However, since UniMODE aims to address both indoor and outdoor 3D object detection using a unified model structure and network weight, its BEV feature must cover a large perception area while still utilizing small BEV grids, which poses a massive challenge due to the limited GPU memory.
To overcome this challenge, we propose a solution that involves partitioning the BEV space into uneven grids, in contrast to the even grids utilized by existing detectors. As depicted in the bottom part of Fig. 2, we achieve this by employing a smaller size of grids closer to the camera and larger grids for those farther away. This approach enables UniMODE to effectively perceive a wide range of objects while maintaining small grid sizes for objects in close proximity. Importantly, this does not increase the total number of grids, thereby avoiding any additional computational burden. Specifically, assuming there are $N_{z}$ grids in the depth axis and the depth range is $(z_{min},z_{max})$, the grid size of the $i_{\rm th}$ grid $z_{i}$ is set to:
@@
Notably, the mathematical form in Eq. 1 is similar to the linear-increasing discretization of depth bin in CaDDN [26], while the essence is fundamentally different. In CaDDN, the feature projection distribution is adjusted to allocate more features to grids closer to the camera. In experiments, we observe that this adjustment results in a more imbalanced BEV feature, i.e., denser features in closer grids and more empty grids in farther grids. Since features in all grids are extracted by the same network, this imbalance degrades the performance. By contrast, our uneven BEV grid approach enhances detection precision by making the feature density more balanced.
3.4 Sparse BEV Feature Projection
The step of transforming the camera view feature into the BEV space is quite computationally expensive due to its numerous projection points. Specifically, considering the image feature $F_{i}\in\mathbb{R}^{C_{i}\times 1\times H_{f}\times W_{f}}$ and depth feature $F_{d}\in\mathbb{R}^{1\times C_{d}\times H_{f}\times W_{f}}$, the projection feature $F_{p}\in\mathbb{R}^{C_{i}\times C_{d}\times H_{f}\times W_{f}}$ is obtained by multiplying $F_{i}$ and $F_{d}$. Therefore, the projection point number $C_{i}\times C_{d}\times H_{f}\times W_{f}$ increases dramatically as the growth of $C_{d}$. The heavy computational burden of this feature projection step restricts the BEV feature resolution, and thus hinders unifying indoor and outdoor 3D object detection.
In this work, we observe that most projection points in $F_{p}$ are unnecessary because their values are quite tiny. This is essentially because of the small corresponding values in $F_{d}$, which imply that the model predicts there is no target in these specific BEV grids. Hence, the time spent on projecting features to these unconcerned grids can be saved.
Based on the above insights, we propose to remove the unnecessary projection points based on a pre-defined threshold $\tau$. Specifically, we eliminate the projection points in $F_{p}$ whose corresponding depth confidence of $F_{d}$ is smaller than $\tau$. In this way, most projection points are eliminated. For instance, when setting $\tau$ to 0.001, about 82.6% of projection points can be excluded.
3.5 Unified Domain Alignment
Heterogeneous domain distributions exist in diverse scenarios and we address this challenge via feature and loss views.
@@ 3.5 Unified Domain Alignment
Obtaining $(\alpha,\beta)$, we employ them to adjust the distribution of $\hat{x}_{l}^{(b,l,c)}$ with respect to $\bar{x}_{l}^{(b,l,c)}=\alpha\cdot\hat{x}_{l}^{(b,l,c)}+\beta$, where $\bar{x}_{l}^{(b,l,c)}$ denotes the updated distribution. In this way, the feature distribution in UniMODE can be adjusted according to input images self-adaptively, and the increased parameters are negligible. Additionally, when an image unseen in the training set is input, DALN still works well, because the unseen image can still be classified as a weighted combination of these $D$ domains.
@@
The two-stage detection architecture described in Section 3.2 can alleviate the aforementioned problem to some extent, because it helps the detector concentrate on foreground objects, and the unlabeled objects are overlooked to compute loss. To address this problem further, we devise a simple strategy, i.e., the class alignment loss. Specifically, denoting the label space of the $i_{\rm th}$ dataset as $\Omega_{i}$, we compute loss on the $i_{\rm th}$ dataset as:
@@
where $l(\cdot)$, $y$, $\bar{y}$, $\mathcal{B}$ are the loss function, class prediction, class label, and background class, respectively. $\gamma$ is a factor for reducing the punishment to classes not included in the label space of this sample.
@@
[…]
@@
Table 2: Detailed performance of UniMODE on various sub-datasets in Omni3D. The detectors are trained and evaluated using the whole Omni3D training and testing data. The results of adopting two different backbones are presented.
4 Experiment
Implementation details. The perception ranges in the X-axis, Y-axis, and Z-axis of the camera coordinate system are $(-30,30)$, $(-40,40)$, $(0,80)$ meters, respectively. If without a special statement, the BEV grid resolution is $(60,80)$. The factor $\gamma$ defined in the class alignment loss is set to 0.2. $M$ and $N$ are set to 100. The adopted optimizer is AdamW, and the learning rate is set to $12e^{-4}$ for a batch size of 192. The experiments are primarily conducted on 4 A100 GPUs.
@@ 4 Experiment
Dataset. The experiments in this section are performed on Omni3D [5], the sole large-scale 3D object detection benchmark encompassing both indoor and outdoor scenes. Omni3D is built upon six well-known datasets including KITTI [8], SUN-RGBD [29], ARKitScenes [3], Objectron [1], nuScenes [6], and Hypersim [27]. Among these datasets, KITTI and nuScenes focus on urban driving scenes, which are real-world outdoor scenarios. SUN-RGBD, ARKitScenes, and Objectron primarily pertain to real-world indoor environments. Compared with outdoor datasets, the required perception ranges of indoor datasets are smaller, and the object categories are more diverse. Hypersim, distinct from the aforementioned five datasets, is a virtually synthesized dataset. Thus, Hypersim allows for the annotation of object classes that are challenging to label in real scenes, such as transparent objects (e.g., windows) and very thin objects (e.g., carpets). The Omni3D dataset comprises a total of 98 object categories and 3 million 3D box annotations, spanning 234,000 images. The evaluation metric is $\rm AP_{3D}$, which reflects the 3D Intersection over Union (IoU) between 3D box prediction and label.
@@
Experimental settings. As Omni3D is a large-scale dataset, training models on it necessitates many GPUs. For example, the authors of Cube RCNN run each experiment with 48 V100s for 4$\sim$5 days. In this work, the experiments in Section 4.1 are performed in the high computing resource setting (the input image resolution is $1280\times 1024$, the backbone is ConvNext-Base [20], all training data). Since our computing resources are limited, unless explicitly stated otherwise, the remaining experiments are conducted in the low computing resource setting (the input resolution is $640\times 512$, the backbone is DLA34 [37], fixed 20% of all training data sampled from all 6 sub-datasets).
4.1 Performance Comparison
In this part, we compare the performance of the proposed detector with previous methods. Among them, Cube RCNN is the sole detector that also explores unified detection. BEVFormer [15] and PETR [18] are two popular BEV detectors, and we reimplement them in the Omni3D benchmark to get the detection scores. The performance of the other compared detectors is obtained from [5]. All the results are given in Table 1. In addition, we present the detailed detection scores of UniMODE on various sub-datasets in Omni3D as Table 2.

### v5

@@ III-C Uneven BEV Grid
Therefore, the BEV grid sizes of current outdoor detectors are typically large to accommodate the vast perception range, while those of indoor detectors are small because of the intricate indoor scenes. However, since UniMODE aims to address both indoor and outdoor 3D object detection using a unified model structure and network weight, its BEV feature must cover a large perception area while still utilizing small BEV grids, which poses a massive challenge due to the limited GPU memory. To overcome this challenge, we propose a solution that involves partitioning the BEV space into uneven grids, in contrast to the even grids utilized by existing detectors. As depicted in the bottom part of Fig. 2, we achieve this by employing a smaller size of grids closer to the camera and larger grids for those farther away. This approach enables UniMODE to effectively perceive a wide range of objects while maintaining small grid sizes for objects in close proximity. Importantly, this does not increase the total number of grids, thereby avoiding any additional computational burden. Specifically, assuming there are $N_{z}$ grids in the depth axis and the depth range is $(z_{min},z_{max})$, the grid size of the $i_{\rm th}$ grid $z_{i}$ is set to:
@@
Notably, the mathematical form in Eq. 1 is similar to the linear-increasing discretization of depth bin in CaDDN [22], while the essence is fundamentally different. In CaDDN, the feature projection distribution is adjusted to allocate more features to the grids that are closer to the camera. In experiments, we observe that this adjustment results in a more imbalanced BEV feature, i.e., denser features in closer grids and more empty grids in farther grids. Since features in all grids are extracted by the same network, this imbalance degrades the 3D object detection performance. By contrast, our uneven BEV grid approach enhances detection precision by making the feature density more balanced.
III-D Sparse BEV Feature Projection
The step of transforming the camera view feature into the BEV space is quite computationally expensive due to its numerous projection points. Specifically, considering the image feature $F_{i}\in\mathbb{R}^{C_{i}\times 1\times H_{f}\times W_{f}}$ and depth feature $F_{d}\in\mathbb{R}^{1\times C_{d}\times H_{f}\times W_{f}}$, the projection feature $F_{p}\in\mathbb{R}^{C_{i}\times C_{d}\times H_{f}\times W_{f}}$ is obtained by multiplying $F_{i}$ and $F_{d}$. Therefore, the projection point number $C_{i}\times C_{d}\times H_{f}\times W_{f}$ increases dramatically as the growth of $C_{d}$. The heavy computational burden of this feature projection step restricts the BEV feature resolution, and thus hinders unifying indoor and outdoor 3D object detection. In this work, we observe that most projection points in $F_{p}$ are unnecessary because their values are quite tiny. This is essentially because of the small corresponding values in $F_{d}$, which imply that the model predicts there is no target in these specific BEV grids. Hence, the time spent on projecting features to these unconcerned grids can be saved.
Based on the above insights, we propose to remove the unnecessary projection points based on a pre-defined threshold $\tau$. Specifically, we eliminate the projection points in $F_{p}$ whose corresponding depth confidence of $F_{d}$ is smaller than the threshold $\tau$. In this way, most projection points are eliminated. For instance, when setting $\tau$ to 0.001, about 82.6% of projection points can be excluded.
III-E Unified Domain Alignment
Heterogeneous domain distributions exist in diverse scenarios and we address this challenge from the feature view and loss view separately.
@@ III-E Unified Domain Alignment
Obtaining $(\alpha,\beta)$, we employ them to adjust the distribution of $\hat{x}_{l}^{(b,l,c)}$ with respect to $\bar{x}_{l}^{(b,l,c)}=\alpha\cdot\hat{x}_{l}^{(b,l,c)}+\beta$, where $\bar{x}_{l}^{(b,l,c)}$ denotes the updated distribution. In this way, the feature distribution in UniMODE can be adjusted according to the input images self-adaptively, and the increased learnable parameters are negligible. Additionally, when an image unseen in the training set is input to the detector, DALN still works well, because the unseen image can still be classified as a weighted combination of these $D$ domains.
@@
The two-stage detection architecture described in Section III-B can alleviate the aforementioned problem to some extent, because it helps the detector concentrate on foreground objects, and the unlabeled objects are overlooked to compute loss. To address this problem further, we devise a simple strategy, i.e., the class alignment loss. Specifically, denoting the label space of the $i_{\rm th}$ dataset as $\Omega_{i}$, we compute loss on the $i_{\rm th}$ dataset as:
@@
where $l(\cdot)$, $y$, $\bar{y}$, $\mathcal{B}$ are the loss function, class prediction, class label, and background class, respectively. $\gamma$ is a factor for reducing the punishment to the classes that are not included in the label space of this sample.
IV Unified Multi-modal Detection
In this section, we describe how we achieve unified multi-modal 3D object detection. Specifically, due to the lack of a proper benchmark, we first explain how we construct the unified multi-modal 3D object detection benchmark MM-Omni3D in Section IV-A. Then, we present the data statistics of the data in this benchmark in Section IV-B. In Section IV-C, we elaborate on the overall framework of the developed multi-modal detector MM-UniMODE, which is extended from the aforementioned UniMODE. Finally, the designed multi-modal feature fusion strategy MIC is introduced in Section IV-D.
IV-A MM-Omni3D Benchmark
MM-Omni3D consists of six datasets, i.e., SUN-RGBD [38], ARKitScenes [11], Objectron [41], Hypersim [42], KITTI [39], and nuScenes [43]. Among them, SUN-RGBD, ARKitScenes, Objectron, and Hypersim are indoor datasets, while KITTI and nuScenes are outdoor datasets. Besides, Hypersim is a synthesized dataset and the other five datasets are collected from real scenes by different sensors.
A multi-modal dataset needs to incorporate at least one depth sensor in addition to a camera. This depth sensor can be of various types, including lidar, radar, RGBD camera, time-of-flight camera, structure from motion, etc. Previous datasets usually use the same depth sensor across the whole dataset [44]. Therefore, the depth point distributions in these datasets are highly similar, and the trained detectors often only perform well when they are tested in scenes that employ the same type of depth sensor. To satisfy the diverse requirements of practical applications, we argue that a multi-modal 3D object detector should utilize camera images and point clouds from various kinds of depth sensors. This is because training a high-performance detector consumes a lot of data. However, collecting multi-modal data is expensive. It is often impractical to prepare abundant multi-modal training data in every specific scene. To handle this challenge, training a universal detector by combining data from various available scenes is feasible.
Figure 5: (a) Camera-view image. (b) Camera-view depth map. (c)$\sim$(f) The processed point clouds observed from the front view, top view, back view, and side view, respectively. As shown, unlike previous 3D object detection datasets [38] that store the point cloud of the whole scene, MM-Omni3D only provides the points visible from the camera view, which better meets the practical situation of online multi-modal 3D object detection.
Nevertheless, the depth from various sensors differs not only in data formats but also in the employed 3D coordinate systems. For example, the depth could be a depth map [39] as shown in Fig. 5 (b) or point cloud [43] in Fig. 5 (c). To address this problem, we transform all depth information into a unified point cloud format and the same camera-view coordinate system. Given that 3D object detection is an online perception task, the multi-modal detector should only have access to depth points that are directly observable by the perception sensors. However, in some data samples, the point cloud is collected in an offline manner, meaning the point cloud for the entire scene is gathered [38]. To solve this problem, as illustrated in Fig. 5 (c)$\sim$(f), we remove all depth points invisible to the camera view at the current timestamp and only retain the visible points. These visible point clouds are associated with image pixels based on camera parameters in a unified format. After investing significant effort in unifying the data formats of various datasets and evaluation procedures, we obtain the MM-Omni3D benchmark. The evaluation metrics used in this benchmark are the same as Omni3D [10].
Figure 6: The data sample numbers of the train, validation, and test sets in the six datasets incorporated in MM-Omni3D.
IV-B MM-Omni3D Data Statistics
The MM-Omni3D dataset consists of a total of 234,152 data samples and is divided into 6 subsets, each corresponding to a different data scenario as shown in Fig. 1 (e)$\sim$(j). Among these subsets, four are dedicated to indoor object detection, while the remaining two pertain to outdoor scenes. Each subset is further divided into training, validation, and testing sets. The number of samples in these subsets is illustrated as Fig. 6, Specifically, for the entire MM-Omni3D dataset, the data sample numbers of the training, validation, and testing sets are about 175.6K, 19.1K, and 39.5K, respectively. In MM-Omni3D, a total of 50 object classes are annotated. It is important to note that only valid 3D box annotations are considered, where the projected 3D centers are visible in the camera images. In addition, the distribution of instances across classes presents a long tail, indicating a significant imbalance in the instance numbers of different classes. For example, there are 152K Car instances in MM-Omni3D while only 21 Floor mat instances, which suggests that the head class with the most instances is about 7.2K times more than the tail class.
The adopted depth sensors are also diverse in MM-Omni3D. As described in Fig. 1 (f)$\sim$(g), although both KITTI and nuScenes employ lidars, their ray numbers are different, resulting in significant variations in point densities. Similarly, the depth points in Objectron and ARKitScenes are obtained through AR sessions, but their densities are quite different due to the discrepancies in collecting data procedures and post-processing operations. In contrast to the aforementioned datasets, the depth points in SUN-RGBD are collected by multiple types of RGB-D cameras. To the best of our knowledge, MM-Omni3D is the first dataset to include such a wide range of depth sensors. The position distributions of targets in MM-Omni3D are also diverse. It can be observed from Fig. 1 (e)$\sim$(j) that indoor targets are mostly in close proximity to the camera, while outdoor objects are far away. This distribution difference could lead to optimization conflict and training instability.
Figure 7: The overall detection framework of MM-UniMODE. For the clarity of this figure, we omit the details of modules that have been illustrated in Fig. 2. The newly added modules include the point cloud feature extraction branch and the MIC for image-point feature fusion.
IV-C MM-UniMODE Overall Framework
In this part, we elaborate on the overall framework of the unified multi-modal 3D object detector MM-UniMODE, which is extended from the aforementioned detector UniMODE. The pipeline of MM-UniMODE is illustrated as Fig. 7. As shown, the detector takes an image and a point cloud as input. The image is captured by a camera, and the point cloud can be collected by any kind of depth sensor. An image branch and a point branch are constructed to extract features from the image and point cloud separately.
The image branch of MM-UniMODE is the same as UniMODE, and its output includes the image BEV feature, domain confidence, and $M+N$ queries. For the point branch, the input point cloud is first voxelized into 3D voxels and then processed by a pillar encoder employed in SECOND [45]. Next, after the pillar scatter operation that transforms the obtained feature as a BEV feature, a point network (including a point backbone and a point neck) is employed to generate the point BEV feature. Obtaining the image and point BEV feature, we fuse them with the proposed MIC module, which enables the image BEV feature and point BEV feature to complement the information of each other, and then the multi-modal BEV feature is derived. Subsequently, 1 encoder and 6 decoders with the same structure implementation as UniMODE are employed to generate detection results based on the multi-modal BEV feature.
IV-D Mutual Information Collaboration
For every training sample, MM-Omni3D takes an image and a point cloud as input and produces the image BEV feature $B_{I}$ and point BEV feature $B_{P}$. Due to the data modality difference, $B_{I}$ and $B_{P}$ present significantly different characteristics. For example, as shown in Fig. 3, the indoor points are denser and closer to the ego sensors, while the outdoor points are sparser and farther. Since the detector needs to detect both indoor and outdoor targets, the size of the BEV feature has to cover the biggest required perception range in MM-Omni3D. Consequently, in outdoor scenes, a large part of $B_{P}$ could do not contain any depth point, and the learned feature in these positions is invalid. Meanwhile, for $B_{I}$, since monocular depth estimation is an ill-posed problem, the extracted feature may be projected to a false grid, which results in the ambiguity in $B_{I}$. Therefore, both the image and point branches have their limitations. Thus we propose the MIC module to fuse image-point BEV features and address the limitations of two branches by enabling these two branches to learn from each other collaboratively, the basic concept of which is illustrated Fig. 7.
Specifically, given $B_{I}$ and $B_{P}$, we first conduct information interaction between them to obtain the updated BEV feature $\hat{B}_{I}$ and $\hat{B}_{P}$. The interaction is implemented using two convolutional layers. For the image feature, the main problem is the quality of $\hat{B}_{I}$ is unsatisfactory due to imprecise depth estimation. By contrast, the feature in grids of $B_{P}$ that contain points is informative. Therefore, we first identify which grids in $B_{P}$ have depth points and denote the boolean mask that suggests the positions of these grids as $M_{P}$. Then, the feature of these grids in $B_{P}$ and $\hat{B}_{I}$ can be represented as $B_{P}M_{P}$ and $\hat{B}_{I}M_{P}$, respectively. Then, we use $B_{P}M_{P}$ to provide guidance to $\hat{B}_{I}M_{P}$ by computing a guidance loss as:
$L_{P\rightarrow I}=l_{1}(\mathcal{S}(B_{P}M_{P}),\hat{B}_{I}M_{P}),$ (8)
where $l_{1}$ is the mean absolute error loss and $\mathcal{S}(\cdot)$ denotes stopping gradient.
In the point branch, the limitation is that numerous BEV grids exactly contain no depth point, so the learned feature in these grids is unreliable. By contrast, most grids in $B_{I}$ contain meaningful features because they are generated based on an estimated depth distribution [22]. Therefore, we propose to employ $B_{I}$ to guide $\hat{B}_{P}$. Denoting the boolean mask that indicates the positions of grids with depth probabilities bigger than a hyper-parameter $\epsilon$ as $M_{I}$, the guidance loss is formulated as:
$L_{I\rightarrow P}=l_{1}(\hat{B}_{P}M_{I}(1-M_{P}),\mathcal{S}(B_{I}M_{I}(1-M_{P})),$ (9)
where $M_{I}(1-M_{P})$ exactly suggests the positions of grids where the image feature is informative but the point feature is unreliable. Obtaining $\hat{B}_{I}$ and $\hat{B}_{P}$, we fuse them as the multi-modal BEV feature using a convolutional block. The losses $L_{P\rightarrow I}$ and $L_{I\rightarrow P}$ described in Eq. 8 and Eq. 9 are only utilized during training and not applied to inference.
V Experiment
In this section, we conduct extensive experiments to validate the effectiveness of the proposed modules used in UniMODE and MM-UniMODE. Additionally, we also study their out-of-domain generalization characteristics and the benefits of the developed multi-modal 3D object detection benchmark MM-Omni3D on serving as a pre-training dataset. To explain the experimental details clearly, we first present some basic information about these experiments as follows:
Implementation details. The perception ranges in the X-axis, Y-axis, and Z-axis of the camera coordinate system are $(-30,30)$, $(-40,40)$, $(0,80)$ meters, respectively. If without a special statement, the BEV grid resolution is $(60,80)$. The factor $\gamma$ defined in the class alignment loss is set to 0.2. The parameter $\epsilon$ for the MIC module is $5e^{-4}$. $M$ and $N$ are set to 100. The adopted optimizer is AdamW, and the learning rate is set to $12e^{-4}$ for a batch size of 192. Both UniMODE and MM-UniMODE are optimized for a total of 116k training iterations. The experiments in Table I and Table II are performed using 16 A100 GPUs and the remaining experiments are primarily conducted with 4 A100 GPUs.
TABLE I: Performance comparison between UniMODE and other monocular 3D object detectors.
@@ V Experiment
TABLE II: Detailed performance of UniMODE on sub-datasets of Omni3D.
@@
Dataset. The experiments in this section are primarily performed on Omni3D [10] (a large-scale monocular 3D object detection benchmark encompassing both indoor and outdoor scenes) and MM-Omni3D (our developed multi-modal 3D object detection benchmark introduced in Section IV-A). Both these two benchmarks are built upon six well-known datasets including KITTI [39], SUN-RGBD [38], ARKitScenes [11], Objectron [41], nuScenes [43], and Hypersim [42]. Among these datasets, KITTI and nuScenes focus on urban driving scenes, which are real-world outdoor scenarios. SUN-RGBD, ARKitScenes, and Objectron primarily pertain to real-world indoor environments. Compared with outdoor datasets, the required perception ranges of indoor datasets are smaller, and the object categories are more diverse. Hypersim, distinct from the aforementioned five datasets, is a synthesized dataset. Thus, Hypersim allows for the annotation of object classes that are challenging to label in real scenes, such as transparent objects (e.g., windows) and thin objects (e.g., carpets). The evaluation metric is $\rm AP_{3D}$, which reflects the 3D Intersection over Union (IoU) between 3D box predictions and labels.
@@
Experimental settings. As Omni3D is a large-scale dataset, training models on it necessitates many GPUs. For example, the authors of Cube RCNN run each experiment with 48 V100s for 4$\sim$5 days. In this work, the experiments in Section V-A are performed in the high computing resource setting (the input image resolution is $1280\times 1024$, the backbone is ConvNext-Base [48], all training data). Since our computing resources are limited, unless explicitly stated otherwise, the remaining experiments are conducted in the low computing resource setting (the input resolution is $640\times 512$, the backbone is DLA34 [49] if without a special statement, fixed 20% of all training data sampled from all 6 sub-datasets).
TABLE III: Performance comparison between MM-UniMODE and other multi-modal 3D object detectors.
Detector Backbone ${\rm AP_{3D}^{sun}}\uparrow$ ${\rm AP_{3D}^{hyp}}\uparrow$ ${\rm AP_{3D}^{ark}}\uparrow$ ${\rm AP_{3D}^{obj}}\uparrow$ ${\rm AP_{3D}^{kit}}\uparrow$ ${\rm AP_{3D}^{nus}}\uparrow$ ${\rm AP_{3D}^{in}}\uparrow$ ${\rm AP_{3D}^{out}}\uparrow$ ${\rm AP_{3D}}\uparrow$
CMT [30] Convext-Base 12.3% 5.3% 36.9% 54.9% 15.7% 18.4% 19.7% 22.7% 22.8%
CMT [30] VoVNet 8.0% 4.1% 26.9% 50.5% 13.0% 17.2% 17.4% 18.9% 18.5%
BEVFusion [29] Convext-Base 12.3% 5.6% 35.8% 59.1% 22.3% 26.1% 21.6% 26.4% 24.1%
BEVFusion [29] VoVNet 11.9% 5.4% 35.7% 52.1% 23.9% 35.7% 21.8% 33.0% 25.1%
FusionFormer [50] Convext-Base 13.8% 5.9% 37.2% 53.1% 21.3% 24.9% 23.5% 25.9% 24.7%
FusionFormer [50] VoVNet 13.4% 5.7% 37.3% 53.9% 25.0% 36.6% 24.0% 35.2% 27.0%
UniMODE Convext-Base 15.7% 4.8% 29.2% 49.0% 13.8% 17.6% 17.3% 20.2% 19.4%
MM-UniMODE Convext-Base 17.1% 6.1% 39.2% 57.1% 23.1% 26.2% 24.6% 27.2% 26.1%
MM-UniMODE VoVNet 16.6% 6.3% 39.5% 59.8% 27.1% 39.1% 25.2% 36.5% 28.6%
V-A Monocular Performance Comparison
[…]